# The N3LO Twist-2 Matching of Helicity TMDs and SIDIS $q_{\ast}$ Spectrum

### v3

@@ 1 Introduction
The principle of factorization underlies all theoretical predictions of QCD. To the UV, the renormalization of $\alpha_{s}$ is independent of the external momenta flowing into the Green functions, so that UV physics is completely factorized from collider measurements performed on asymptotic final states. To the IR, long-distance physics is parametrized while short-distance dynamics are computed from Feynman amplitudes, again thanks to the asymptotic freedom. UV renormalization and IR factorization scheme ambiguities drop out in the relations among physical quantities.
@@ 2 The TMD Factorization of $q_{\ast}$ Spectrum
Once the hard and soft-collinear degrees of freedom are disentangled, the next step is to enumerate all allowed hard events, and within each hard label, zoom in to the soft-collinear subprocesses and count over particles in a proper way: the classes of momenta entering into the soft-collinear phase space differ only by their typical size of energy or rapidity, as a result, the phase space should be integrated over with a rapidity cut-off $\nu$, here we employ exponential cut-off scheme Li:2016axz . Additionally, a zero-bin subtraction is performed to remove double-counting between collinear and soft sectors, yielding genuine collinear beam functions Manohar:2006nz . Consequently, we have hard-collinear-soft factorization
@@ 3 Renormalization Group Evolution and $q_{\ast}$ Spectrum with RG Improved Perturbation Theory
The factorization framework of eq. (3) was pioneered by Collins, Soper, and Sterman (CSS) Collins:1981uk ; Collins:1981uw ; Collins:1984kg and expanded in Collins:2011zzd ; Ji:2004wu ; deFlorian:2001zd ; Catani:2000vq ; Catani:2010pd ; Aybat:2011zv ; Ji:2005nu , and later reformulated in the framework of Soft Collinear Effective Theory Becher:2010tm ; Becher:2011xn ; Becher:2012yn ; Echevarria:2011epo ; Echevarria:2012js ; Echevarria:2014rua ; Ebert:2019tvc ; Collins:2017oxh .
@@
The N3LL predictions for eq. (3) is shown in fig. (1,2) for both unpolarized and polarized distributions, where we incorporate the non-perturbative model given in Ref Sun:2014dqm
@@ 4.1 Operator definitions for the TMD helicity PDFs and FFs
where the factor $\mathcal{N}_{\text{sch.}}$ depends on the definition of $\gamma_{5}$ in dimensional regularization. Furthermore, the integral in Eq. (63) is properly defined only with a rapidity cutoff implemented through the exponential rapidity regulator $e^{-b_{0}\tau\frac{P\cdot K}{P^{+}}}$ Li:2016axz ; Luo:2019hmp , which effectively restricts the total energies of the collinear radiations by a rapidity scale $\nu=1/\tau$. The TMD FFs are defined as crossings of TMD beam functions, and according to parton-hardron frame relation given in eq. (20Collins:2011zzd ; Luo:2019hmp ; Luo:2019bmw , we have, in dimensional regularization
@@ 4.4.1 Space-like results in $\overline{\text{MS}}$}
The space-like helicity dependent splitting functions were obtained in Moch:2014sna ; Moch:2015usa ; Blumlein:2021enk ; Blumlein:2021ryt ; Blumlein:2022gpp , by comparison we find all agreement except for the term with cubic color structure $d_{abc}^{2}$ for ‘sea’ quark difference $N_{f}(\Delta P_{qq^{\prime}}-\Delta P_{q\bar{q}^{\prime}})$:  [^1] Our results agree at transcendental weight-4, but discrepancies appear for the weight-3 terms, indicating differences of $\gamma_{5}$ prescriptions between the one used in Ref Moch:2015usa and our tensor reduction prescription following Ravindran:2003gi .
$\Delta P_{d_{abc}^{2}}^{S,(2)}-\Delta P_{d_{abc}^{2}}^{\text{Moch:2015}}=\frac{16N_{f}}{3}\frac{d_{abc}^{2}}{N_{c}}\bigg[-18\zeta_{3}+3H_{0}+84H_{2}+36H_{3}+12H_{-2,0}+(1-z)\times\left(174H_{1}-12H_{1,0,0}-84\zeta_{2}\right)+(1+z)\times\left(36H_{-1,0}-48H_{-1,2}-24H_{-1,0,0}+48H_{-1}\zeta_{2}-36H_{0}\zeta_{2}\right)+z\times\left(-174H_{0}+90H_{2}+12H_{3}-24H_{-2,0}+54H_{0,0}+24H_{0,0,0}-54\zeta_{2}-72\zeta_{3}\right)\bigg]\,.$ (122)

### v4

@@ 1 Introduction
The principle of factorization underlies all theoretical predictions of QCD. To the UV, the renormalization of $\alpha_{s}$ is independent of the external momenta flowing into the Green functions, so that UV physics is completely factorized from collider measurements performed on asymptotic final states. To the IR, long-distance physics is parametrized while short-distance dynamics are computed from Feynman amplitudes, again thanks to the asymptotic freedom. UV and IR cutoff dependence drop out in the relations among physical quantities.
@@ 2 The TMD Factorization of $q_{\ast}$ Spectrum
Once the hard and soft-collinear degrees of freedom are disentangled, the next step is to enumerate all allowed hard events, and within each hard label, zoom in to the soft-collinear subprocesses and count over particles in a proper way: the classes of momenta entering into the soft-collinear phase space differ only by their typical size of energy or rapidity, as a result, the phase space should be integrated over with a rapidity cutoff $\nu$, here we employ exponential cutoff scheme Li:2016axz . Additionally, a zero-bin subtraction is performed to remove double-counting between collinear and soft sectors, yielding genuine collinear beam functions Manohar:2006nz . Consequently, we have hard-collinear-soft factorization
@@ 3 Renormalization Group Evolution and $q_{\ast}$ Spectrum with RG Improved Perturbation Theory
The factorization framework of eq. (3) was pioneered by Collins, Soper, and Sterman (CSS) Collins:1981uk ; Collins:1981uw ; Collins:1984kg and expanded in Collins:2011zzd ; Ji:2004wu ; deFlorian:2001zd ; Catani:2000vq ; Catani:2010pd ; Aybat:2011zv ; Ji:2005nu , and later reformulated in the framework of Soft Collinear Effective Theory Becher:2010tm ; Becher:2011xn ; Becher:2012yn ; Echevarria:2011epo ; Echevarria:2012js ; Echevarria:2014rua ; Ebert:2019tvc ; Collins:2017oxh . The N3LL predictions for eq. (3) is shown in fig. (1,2) for both unpolarized and polarized distributions, where we incorporate the non-perturbative model given in Ref Sun:2014dqm
@@ 4.1 Operator definitions for the TMD helicity PDFs and FFs
where the factor $\mathcal{N}_{\text{sch.}}$ depends on the definition of $\gamma_{5}$ in dimensional regularization. Furthermore, the integral in Eq. (63) requires a rapidity cutoff to be well defined. In this work, we implement such a cutoff via the exponential rapidity regulator $e^{-b_{0}\tau\frac{P\cdot K}{P^{+}}}$ Li:2016axz ; Luo:2019hmp , which effectively restricts the total energies of the collinear radiations by a rapidity scale $\nu=1/\tau$. The TMD FFs are defined as crossings of TMD beam functions, and according to parton-hardron frame relation given in eq. (20Collins:2011zzd ; Luo:2019hmp ; Luo:2019bmw , we have, in dimensional regularization
@@ 4.4.1 Space-like results in $\overline{\text{MS}}$}
The space-like helicity dependent splitting functions were obtained in Moch:2014sna ; Moch:2015usa ; Blumlein:2021enk ; Blumlein:2021ryt ; Blumlein:2022gpp , by comparison we find all agreement except for the term with cubic color structure $d_{abc}^{2}$ for ‘sea’ quark difference $N_{f}(\Delta P_{qq^{\prime}}-\Delta P_{q\bar{q}^{\prime}})$:  [^1] Agreement is found at transcendental weight four, whereas discrepancies arise in the weight-three terms. The author is also grateful to the authors of Ref. Behring:2025avs for pointing out a typo in the $\zeta_{2}$ term in the output of this formula. Once corrected, the expression is in full agreement with the file dPSLarin.m accompanying the present work and with their independent calculations.
$\Delta P_{d_{abc}^{2}}^{S,(2)}-\Delta P_{d_{abc}^{2}}^{\text{Moch:2015}}=\frac{16N_{f}}{3}\frac{d_{abc}^{2}}{N_{c}}\bigg[-84\zeta_{2}-18\zeta_{3}+3H_{0}+84H_{2}+36H_{3}+12H_{-2,0}+(1-z)\times\left(174H_{1}-12H_{1,0,0}\right)+(1+z)\times\left(36H_{-1,0}-48H_{-1,2}-24H_{-1,0,0}+48H_{-1}\zeta_{2}-36H_{0}\zeta_{2}\right)+z\times\left(-174H_{0}+90H_{2}+12H_{3}-24H_{-2,0}+54H_{0,0}+24H_{0,0,0}-54\zeta_{2}-72\zeta_{3}\right)\bigg]\,.$ (122)